# Fixed Point Explainability

### v2

@@ Abstract
This paper introduces a formal notion of fixed point explanations, inspired by the “why regress” principle, to assess, through recursive applications, the stability of the interplay between a model and its explainer. Fixed point explanations satisfy properties like minimality, stability, and faithfulness, revealing hidden model behaviours and explanatory weaknesses. We define convergence conditions for several classes of explainers, from feature-based to mechanistic tools like Sparse AutoEncoders, and we report quantitative and qualitative results. [^1] The code to replicate all the experiments in the paper is available here: https://github.com/EmanueleLM/fixed-point-explainability.
@@ 2 Methodology
Figure 2: A $P$-fixed point LIME explanation for a VGG16 network [72] on CIFAR10 [47]. At each iteration, part of the input is occluded (the yellow patches) by the explainer but the model’s classification is preserved and correct. After iteration $4$ (the fixed point), the correct classification will hold up to infinity.
@@ 2.1.3 Faithful, Stable, and Progressive Explanations
For example, the invariance of the model to local perturbations [57], e.g., in $\mathcal{L}_{p}$-norm, can be expressed as $f(x^{\prime})=y,\forall x^{\prime}\ .\ ||x-x^{\prime}||_{p}\leq\rho$, and enclosed in the set of properties $P$. A similar argument applies to the stability of the explainer. We nonetheless notice that such properties are computationally hard to verify [46].
@@ 3.1 Feature-based Explainers
In other words, for any input, under monotonicity of $(\varepsilon\circ s)$ w.r.t. the $\subseteq$ operator, i.e., $A\subseteq B\implies\varepsilon(s(A))\subseteq\varepsilon(s(B))$, a fixed point always exists.
@@ 3.3 Mechanistic Interpretability and Sparse AutoEncoders
In other words, a $P$-fixed point explanation preserves the next-token prediction of the LLM: the condition $\mathbb{P}(f(x|\ \text{do}(z\leftarrow z_{k})))\approx\mathbb{P}(f(x|\ \text{do}(z\leftarrow z_{n+1})))$ guarantees that the token distribution at each iteration does not change much. While a lossless SAE has, trivially, for any input, a fixed point that converges after one iteration, in Appendix C we expand on the general conditions to guarantee the existence and uniqueness of a fixed point and $P$-fixed point explanation, distinguishing between the linear and non-linear case. In the experiments, we show how popular pre-trained SAEs lack consistency as the features’ reconstruction does not preserve the LLM predicted tokens, posing serious issues regarding trusting their explanations and the “disentangled” features they are supposed to extract.
@@ 4.1 Feature-based Explainers
On three standard explainers in literature, namely LIME, SHAP, and LRP, and three standard datasets, namely MNIST, FashionMNIST, and CIFAR10, we compute fixed point and $P$-fixed point explanations on a VGG16 network. Quantitative results for each dataset and explainer are reported in Table 1. LIME explanations are very short and inconsistent (i.e., the recursive steps do not preserve the class predicted by the model originally): we impute this to LIME being a local explainer. On the other hand, LRP and SHAP explanations are more consistent, as they act globally and are less affected by spurious features in the input. Figure 7 reports $P$-fixed points for each class in a dataset, revealing a minimal subset of the anchoring features of each class.
@@ 4.2 Prototype-based Explainers
Figure 7: Fixed point explanations of MNIST (top), FashionMNIST (centre) and CIFAR10 (bottom) classes, that preserve the correct model’s classification upon convergence and up to infinity. No fixed point was found for the class “dog” in the entire CIFAR10 dataset: a curious case we further discuss in Appendix E.1.
@@ 4.3 SAE
Our LLM of reference is Llama-3.1-8B [32], which we explain with an open-source SAE by EleutherAI [24]. We compute fixed point and $P$-fixed point explanations (Def. 3.2) on a dataset containing three types of questions: numerical, multiple choice, and general knowledge. The dataset has been synthetically generated by GPT4 [63] to cover domains requiring a single-token answer. More details are provided in Appendix E.2. We apply the SAE at the last token position of the last LLM layer ($24$), so that we do not have to consider successive features (merging LLMs features across token positions is an open research question [10]).
Table 2 shows that the SAE iterations converge to a fixed point $100\%$ of the time. The Jaccard similarity between features before and after one recursive SAE iteration is greater than $86\%$ for all our question types, indicating a significant overlap between activating features in consecutive SAE iterations. In less than $100$ iterations and for all the prompts, this overlap reaches $100\%$. However, the low Jaccard similarity between starting features and $P$-fixed point explanation features suggests that the SAE identifies different features at each iteration. Regarding correct answers given by the LLM, the first and the first recursive SAE iterations have similar predictions. However, for $P$-fixed point explanations that number drops to $0\%$ across all question types, indicating that, upon convergence and thus stability, SAEs identify features that diverge from those it initially discovers.
Question Type Setting # iterations Correctness Correctness (Top5) Jaccard
Numerical base 00 $0.5446$ $0.9802$
single $1$ $0.6238$ $0.9505$
double $2$ $0.6139$ $0.9505$ $0.8720$
constant $29\pm 12$ $0.0000$ $0.0000$ $0.0169$
Multiple-Choice base 00 $0.7300$ $0.9700$
single $1$ $0.7600$ $0.9700$
double $2$ $0.7500$ $0.9900$ $0.8653$
constant $17\pm 4$ $0.0000$ $0.0000$ $0.0268$
General-Knowledge base 00 $0.1667$ $0.3529$
single $1$ $0.1863$ $0.3627$
double $2$ $0.1667$ $0.3529$ $0.8831$
constant $31\pm 13$ $0.0000$ $0.0000$ $0.0182$
Table 2: Performance metrics for Llama-3.1-8B as explained by an SAE across three categories with around $100$ questions each. Setting specifies how many SAE iterations are applied to the LLM, i.e., 00, $1$, $2$, or those sufficient to converge (specified in #iterations). Correctness (i.e., the proportion of correct answers) concerns $P$-fixed point explanations for the top token, or across the top five tokens. Jaccard similarity is the proportion of overlapping features with the first SAE iteration.
@@ 5 Related Work
The above mentioned works mainly focus on robust explanations, i.e., those guaranteed to remain invariant for the model. A complementary line of research devises robust explainers, i.e., methods that provide explanations that are robust for the explainer [18, 23, 54, 77, 79]. A few works analyse common explainers and how their output changes when explaining different models and initial conditions [5, 45].
@@ Appendix A Explainers, Endomorphisms, and Support Functions
When an explainer is not an endomorphism, one needs a support function $s$ that, in composition with the explainer, makes the composition $\varepsilon\circ s$ an endomorphism so that the notion of fixed point explanation applies. For prototype-based explainers such as LIME, SHAP and LRP, it is natural to treat the features identified as important as those the support function keep constant for the next iteration. On the other hand, the features that contribute negatively to the classification can be occluded (e.g., set to zero or their value decreased).
Different approaches are viable: explainers such as Anchors [68], identify the features that are responsible for the correct classification while optimising precision and coverage. The left-out features are sampled from a target distribution (e.g., that of an LLM where the left-out features are sampled conditioning on those Anchors identifies as relevant for the explanation).
For prototype-based explanations, as well as SAEs, the support function is not required as the explainability process is an endomorphism. As we discuss in the main paper and in Appendix D.1 and D.2, rule-based explainers and LLMs require a careful characterisation of the support function, as characterising them as monotonic explainers misses the core characteristics of the models we would like to explain. We thus leave their characterisation, as well as a theoretical extension of our framework, as future work.
@@ C.1 On Contractive SAE and fixed point explanations
A linear SAE has the following mathematical formulation: $\varepsilon(z;f)=zW^{E}W^{D}$, which we refer to as $zW$ for the linearity of the transformations involved. A condition to have a fixed point in a recursive SAE is that the SAE transformation $W=W^{E}W^{D}$ is “contractive”. In other words, that is equivalent to checking whether (i) $W$ is diagonalisable and (ii) its spectral norm is lower or strictly lower than $1$. [^3] The spectral norm of a diagonalisable projection $W=A\Sigma A^{T}$ is the square root of its maximum eigenvalue. In the general case of a non-squared matrix, one would look at the Lipschitz of the transformations (easy for the linear case and bounded by the norm of the matrix), or at other techniques such as SVD.
Suppose all the eigenvalues of $W$ are at most $1$. In that case, the repetition of the application of the SAE will converge to zero for all the dimensions that correspond to the eigenvectors with eigenvalues $<1$. In contrast, they will converge to a non-zero fixed point for those equal to $1$.
Final note. If there is one single eigenvalue with value $1$, and (i) and (ii) hold, then the fixed point is unique and the same for any input; otherwise, it is different for each input (that’s what we expect in general by a non-trivial linear SAE). If an eigenvalue is complex, we look at its norm: if it’s complex but its norm is $<1$, the process goes to zero (in a spiral and via rotations). If it equals $1$, it rotates forever and does not converge. If it is $>1$, it diverges.
@@
A non-linear SAE has the following mathematical formulation: $\varepsilon(z;f)=a(zW^{E})W^{D}$, with $a$ the binary top-$k$ function. Proving that an input has a fixed point requires testing that any combination of the possible $\binom{|h|}{k}$ binary vectors the SAE produces has a fixed point, i.e., that the decoder is contractive. While that requires a combinatorial number of steps, there are a few alternative approaches to mitigate the hardness of the problem. To guarantee the global convergence of any input, one can estimate the Lipschitz constant of the SAE; the $inf$ and $sup$ of each SAE input dimension $z^{(i)}$ give a bound of the Lipschitz constant of the network, from which one can derive the max and min of each output neuron analytically. While a global Lipschitz bound is hard to estimate, or too loose in practice, one can assume the Hamming distance between a hidden representations $h$ and its recursive iteration $h^{\prime}=\text{top-}k(hW^{D}W^{E})$ being bounded (e.g., only a few neurons of the top-$k$ may change between the two).
[…]
@@ D.1 Rule-based Explainers
Let’s consider a model $f:X\xrightarrow{}Y,\ |Y|<\infty$. An explainer is defined as $\varepsilon:X\xrightarrow{}R,\ |R|<\infty$. Suppose $X\in\{0,1\}^{d}$ and $R=\{x^{(i)}=\{0,1,?\}\}$, where $x^{(i)}$ is the i-th feature of $x$ and $\{0,1,?\}$ means that its value is either assigned to 00 or $1$ or undefined. For example, a valid assignment for a bi-dimensional input is $\{0,?\}$, which corresponds to the rule ‘$\text{if}\ x^{(1)}=1\ \ \text{then}\ \ y$’. Given a support function $s$ (e.g., all the points undefined in $R$ are replaced by the mode of that variable in the training data), to guarantee convergence one has to prove, or construct $\varepsilon$ and $s$ such that $(\varepsilon\circ s)(x)$ is, for any perturbation $s$ can induce for undefined variables in an assignment $r$, a stable attractor.
@@ E.1 Feature-based Explainers
Figure 10: For our VGG16 model, there is no $P$ fixed point explanation for the class “dog”. Each successive iteration of the explainer is either classified as “bird”, “cat”, or “horse”.
@@
[…]
The curious case of the missing dog.
As mentioned in Figure 7, we couldn’t find a $P$-fixed point explanation for the class “dog” in the entire CIFAR10 test set (neither with LIME, SHAP, nor LRP). Our empirical results suggest that every successive iteration of the explainer makes the model classify the image as a “cat” ($57\%$ of the time), a “horse” ($27\%$), or a bird ($14\%$). This issue is only partially testified by the accuracy-per-class of the model, which is around $88-90\%$ for each class, but it drops to $59.5\%$ for “dog”. In fact, just by inspecting the accuracy-per-class, one wouldn’t figure out the features the model and the explainer interpret as belonging to the other classes, which our technique identifies, as per Figure 10.

### v3

@@ Abstract
This paper introduces a formal notion of fixed point explanations, inspired by the “why regress” principle, to assess, through recursive applications, the stability of the interplay between a model and its explainer. Fixed point explanations satisfy properties like minimality, stability, and faithfulness, revealing hidden model behaviours and explanatory weaknesses. We define convergence conditions for several classes of explainers, from feature-based to mechanistic tools like Sparse AutoEncoders, and we report quantitative and qualitative results for several datasets and models, including LLMs such as Llama-3.3-70B. [^1] The code to replicate all the experiments in the paper is available here: https://github.com/EmanueleLM/fixed-point-explainability.
@@ 2 Methodology
Figure 2: A $P$-fixed point LIME explanation for a VGG16 network [72] on CIFAR10 [46]. At each iteration, part of the input is occluded (the yellow patches) by the explainer but the model’s classification is preserved and correct. After iteration 44 (the fixed point), the correct classification will hold up to infinity.
@@ 2.1.3 Faithful, Stable, and Progressive Explanations
[…]
@@ 3.1 Feature-based Explainers
In other words, for any input, under monotonicity of $(\varepsilon\circ s)$ w.r.t. the $\subseteq$ operator, i.e., $A\subseteq B\implies\varepsilon(s(A))\subseteq\varepsilon(s(B))$, a fixed point always exists (note that even non-monotonic explainers will behave monotonically when combined with e.g. masking-based support functions, ensuring convergence). We conjecture that similar guarantees may hold even without the monotonicity assumption, which we will pursue as a future direction.
@@ 3.3 Mechanistic Interpretability and Sparse AutoEncoders
In other words, a $P$-fixed point explanation preserves the next-token prediction of the LLM: the condition $\mathbb{P}(f(x|\ \text{do}(z\leftarrow z_{k})))\approx\mathbb{P}(f(x|\ \text{do}(z\leftarrow z_{n+1})))$ guarantees that the token distribution at each iteration changes minimally. While a lossless SAE has, trivially, for any input, a fixed point that converges after one iteration, in Appendix C we expand on the general conditions to guarantee the existence and uniqueness of a fixed point and $P$-fixed point explanation, distinguishing between the linear and non-linear case. In the experiments, we show how popular pre-trained SAEs lack consistency as the features’ reconstruction does not preserve the LLM predicted tokens, posing serious issues regarding trusting their explanations and the “disentangled” features they are supposed to extract.
@@ 4.1 Feature-based Explainers
On three standard explainers in literature, namely LIME, SHAP, and LRP, and three standard datasets, namely MNIST, FashionMNIST, and CIFAR10, we compute fixed point and $P$-fixed point explanations on a VGG16 network. Quantitative results for each dataset and explainer are reported in Table 1. LIME explanations are very short and inconsistent (i.e., the recursive steps do not preserve the class predicted by the model originally): we impute this to LIME being a local explainer. On the other hand, LRP and SHAP explanations are more consistent, as they act globally and are less affected by spurious features in the input. Figure 7 reports $P$-fixed points for each class in a dataset, revealing a minimal subset of the anchoring features of each class. Further results for different support functions are provided in Appendix A.
@@ 4.3 SAE
We focus on Llama-3.3-70b and Llama-3.1-8B [31], which we explain with a open-source SAEs [59, 23]. We compute fixed point and $P$-fixed point explanations (Def. 3.2) on a dataset containing three types of questions: numerical, multiple choice, and general knowledge. The datasets has been synthetically generated by GPT4 [63] to cover domains requiring a single-token answer. More details are provided in Appendix E.2. We apply the SAEs in the deeper layers of LLMs, where the skip connections of transformer layers accumulate features across the model.
Table 2 shows that the SAEs converge to a fixed point $100\%$ of the time, and do so faster and towards more informative fixed points across larger models, hinting at a scaling law for stability and recursive explanation quality. In particular, we see that consecutive SAE iterations of Llama 70b have lower initial Jaccard similarities, converging faster to the FPE (defined as total overlap), and lower KL divergence with the initial logits, which yield $93\%$ P-FPE in numerical questions. In Appendix C.2, we further report results for Llama-3.1-8B.
Figure 7: Fixed point explanations of MNIST (top), FashionMNIST (centre) and CIFAR10 (bottom) classes, that preserve the correct model’s classification upon convergence and up to infinity.
Question Type Setting # Iterations Correctness Correctness (Top5) Jaccard KL Div
Numerical base 0 $0.6436$ $0.9703$ - -
single 11 $0.4158$ $0.7525$ - $0.6665$
double 22 $0.3069$ $0.5941$ $0.3264$ $1.0862$
constant $7\pm 3$ $0.0000$ $0.0000$ $0.0020$ $5.6397$
Multiple-Choice base 0 $0.9700$ $0.9900$ - -
single 11 $0.9700$ $0.9900$ - $0.1482$
double 22 $0.9700$ $0.9900$ $0.2786$ $0.2242$
constant $7\pm 1$ $0.9300$ $0.9900$ $0.0022$ $0.6239$
General-Knowledge base 0 $0.2647$ $0.4706$ - -
single 11 $0.0686$ $0.2059$ - $0.9323$
double 22 $0.0098$ $0.1275$ $0.3413$ $1.7299$
constant $13\pm 12$ $0.0196$ $0.0980$ $0.0019$ $5.3277$
Table 2: Metrics for FPEs of Llama-3.3-70B produced by an SAE across three categories of questions. Setting specifies how many SAE iterations are applied to the LLM, i.e., 0, 11, 22, or the #Iterations sufficient to get to the constant point. Correctness (the proportion of correct answers for the top token, or the top five tokens) is our $P$-fixed point explanation in the constant case. Jaccard similarity is the proportion of overlapping features with the first SAE iteration. The KL divergence is between original logits and those patching the SAE reconstruction in the residual stream.
@@ 5 Related Work
The aforementioned works mainly focus on robust explanations, i.e., those guaranteed to remain invariant for the model. A complementary line of research devises robust explainers, i.e., methods that provide explanations that are robust for the explainer [17, 22, 53, 77, 79]. A few works analyse common explainers and how their output changes when explaining different models and initial conditions [5, 44].
@@ Appendix A Explainers, Endomorphisms, and Support Functions
When an explainer is not an endomorphism, one needs a support function $s$ that, in composition with the explainer, makes the composition $\varepsilon\circ s$ an endomorphism so that the notion of fixed point explanation applies. For prototype-based explanations, as well as SAEs, the support function is not required as the explainability process is an endomorphism. As we discuss in the main paper and in Appendix D.1 and D.2, rule-based explainers and LLMs require a careful characterisation of the support function, as characterising them as monotonic explainers misses the core characteristics of the models we would like to explain. We thus leave their characterisation, as well as a theoretical extension of our framework, as future work.
In contrast, for feature-based explainers, designing good support functions is inherently challenging: a well-designed support must effectively suppress discarded features without excessively distorting the input or reintroducing discarded features (which would disrupt the recursive filtering process). To better understand the sensitivity of our framework to this choice, we conducted additional experiments exploring alternative support functions for feature-based explanations on CIFAR-10. More detailedly, we compared the standard zero-out strategy against three alternatives designed to balance feature suppression and visual coherence:
  • Exponential Decay: masked pixel values are multiplied by a decay factor (0.7), progressively fading discarded regions.
  • Class-wise Average Filling: masked pixels are replaced by class-specific average images, providing a more semantically plausible background.
  • Gaussian Blur: masked pixels are replaced by a locally blurred version of the original image, which preserves low-frequency structure while removing fine details.
We focus the analysis on the LRP explainer, due to its computational efficiency and higher flexibility in the explanation masking. Quantitative results are shown in Table 3. As it can be observed, similar consistency levels are obtained across all methods (differences within approximately 2%-12%) while being less disruptive visually (e.g., Gaussian blurring), demonstrating that our results are robust to the choice of the support function. Figure 8 provides qualitative comparisons, illustrating how each support method influences the appearance of recursive inputs throughout the process.
Metric Zero-out Exp. Decay Class-wise Averaging Gaussian Blur
Self-consistency $\% 2626\%$ $14\%$ $24\%$
Convergence Steps $5.34\pm 0.92$ $69.48\pm 3.42$ $92\pm 1.20 6.6$$64\pm 1.34$
Table 3: Self-consistency rates (%) and convergence steps of LRP on CIFAR-10, under different support functions: zero-out masking, exponential decay (with a factor of 0.7), class-wise average filling and Gaussian blur.
Figure 8: Visual comparison of different support functions for CIFAR10 and LRP. Each row shows a different support method: zero-out, exponential decay, class-wise averaging, and Gaussian blur.
@@ C.1 On Contractive SAE and fixed point explanations
A linear SAE has the following mathematical formulation: $\varepsilon(z;f)=zW^{E}W^{D}$, which we refer to as $zW$ for the linearity of the transformations involved. A condition to have a fixed point in a recursive SAE is that the SAE transformation $W=W^{E}W^{D}$ is “contractive”. In other words, that is equivalent to checking whether (i) $W$ is diagonalisable and (ii) its spectral norm is lower or strictly lower than 11.333The spectral norm of a diagonalisable projection $W=A\Sigma A^{T}$ is the square root of its maximum eigenvalue. In the general case of a non-squared matrix, one would look at the Lipschitz of the transformations (easy for the linear case and bounded by the norm of the matrix), or at other techniques such as SVD.
Suppose all the eigenvalues of $W$ are at most $1$. In that case, the repetition of the application of the SAE will converge to zero for all the dimensions that correspond to the eigenvectors with eigenvalues <1<1. In contrast, they will converge to a non-zero fixed point for those equal to $1$.
Final note. If there is one single eigenvalue with value $1$, and (i) and (ii) hold, then the fixed point is unique and the same for any input; otherwise, it is different for each input (that’s what we expect in general by a non-trivial linear SAE). If an eigenvalue is complex, we look at its norm: if it’s complex but its norm is <1<1, the process goes to zero (in a spiral and via rotations). If it equals $1$, it rotates forever and does not converge. If it is >1>1, it diverges.
@@
A non-linear SAE has the following mathematical formulation: $\varepsilon(z;f)=a(zW^{E})W^{D}$, with $a$ the binary top-$k$ function. Proving that an input has a fixed point requires testing that any combination of the possible $\binom{|h|}{k}$ binary vectors the SAE produces has a fixed point, i.e., that the decoder is contractive. While that requires a combinatorial number of steps, there are a few alternative approaches to mitigate the hardness of the problem. To guarantee the global convergence of any input, one can estimate the Lipschitz constant of the SAE; the infinf and supsup of each SAE input dimension $z^{(i)}$ give a bound of the Lipschitz constant of the network, from which one can derive the max and min of each output neuron analytically. While a global Lipschitz bound is hard to estimate, or too loose in practice, one can assume the Hamming distance between a hidden representations $h$ and its recursive iteration $h^{\prime}=\text{top-}k(hW^{D}W^{E})$ being bounded (e.g., only a few neurons of the top-$k$ may change between the two).
Figure 10: Convergence/divergence of the recursive process of a linear SAE $\varepsilon(x)=xW$. Left: $1000$ random $10\text{x}10$ matrices $W$ that, by construction, have their eigenvalues with norm at most $1$. Right: $1000$ random $10\text{x}10$ matrices that, by construction, have some of their eigenvalues with norm larger 11.
C.2 Additional Experiments with SAE
In Table 5, we report the results for the fixed point explanations we obtained with Llama-3.1-8B. For reference, results for Llama-3.3-70B are reported in the main paper, Table 2.
Question Type Setting # Iterations Correctness Correctness (Top5) Jaccard KL Div
Numerical base 0 $0.5446$ $0.9802$ - -
single 11 $0.6238$ $0.9505$ - $0.1517$
double 22 $0.6139$ $0.9505$ $0.8720$ $0.1910$
constant $29\pm 12$ $0.0000$ $0.0000$ $0.0169$ $10.0825$
Multiple-Choice base 0 $0.7300$ $0.9700$ - -
single 11 $0.7600$ $0.9700$ - $0.1071$
double 22 $0.7500$ $0.9900$ $0.8653$ $0.1680$
constant $17\pm 4$ $0.0000$ $0.0000$ $0.0268$ $9.7767$
General-Knowledge base 0 $0.1667$ $0.3529$ - -
single 11 $0.1863$ $0.3627$ - $0.0819$
double 22 $0.1667$ $0.3529$ $0.8831$ $0.1145$
constant $31\pm 13$ $0.0000$ $0.0000$ $0.0182$ $8.7088$
Table 5: MMetrics for FPEs of Llama-3.1-8B produced by an SAE across three categories of questions. Setting specifies how many SAE iterations are applied to the LLM, i.e., 0, 11, 22, or the #Iterations sufficient to get to the constant point. Correctness (the proportion of correct answers for the top token, or the top five tokens) is our $P$-fixed point explanation in the constant case. Jaccard similarity is the proportion of overlapping features with the first SAE iteration. The KL divergence is between original logits and those patching the SAE reconstruction in the residual stream.
@@ D.1 Rule-based Explainers
Let’s consider a model $f:X\xrightarrow{}Y,\ |Y|<\infty$. An explainer is defined as $\varepsilon:X\xrightarrow{}R,\ |R|<\infty$. Suppose $X\in\{0,1\}^{d}$ and $R=\{x^{(i)}=\{0,1,?\}\}$, where $x^{(i)}$ is the i-th feature of $x$ and $\{0,1,?\}$ means that its value is either assigned to 0 or $1$ or undefined. For example, a valid assignment for a bi-dimensional input is $\{0,?\}$, which corresponds to the rule $\text{if}\ x^{(1)}=1\ \ \text{then}\ \ y$’. Given a support function $s$ (e.g., all the points undefined in $R$ are replaced by the mode of that variable in the training data), to guarantee convergence one has to prove, or construct $\varepsilon$ and $s$ such that $(\varepsilon\circ s)(x)$ is, for any perturbation $s$ can induce for undefined variables in an assignment $r$, a stable attractor.
@@ E.1 Feature-based Explainers
For each dataset, a VGG16 model is trained for $5$ epochs, with an accuracy of $0.94$ on MNIST, $0.85$ on FashionMNIST, and $0.88$ on CIFAR10. MNIST and FashionMNIST images are scaled from $28\text{x}28$ to $32\text{x}32$ pixels to be compatible with the VGG16 network by padding their border.